A baker purchased 10 lb of wheat flour and 11 lb of rye flour for a total cost of $13.60. A second purchase, at the same prices, included 13 lb of wheat flour and 14 lb of rye flour. The cost of the second purchase was $17.50. Find the cost per pound of the wheat flour and of the rye flour.
Answer #1

cost per pound of the wheat flour =$x

cost per pound of the rye flour =$y

Set up equation as

10x+11y=13.60

13x+14y=17.50

From first equation,

11y= 13.6 -10x

y =(13.6 -10x)/11

Substitute this in second equation

13x+14*(13.6 -10x)/11=17.50

Multiply both sides by 11

143x+190.4-140x=192.5

3x=192.5-190.4

3x=2.1

x=2.1/3

x=0.7

Substitute this in y =(13.6 -10x)/11

y =(13.6 -10*0.7)/11 =0.6

hence

cost per pound of the wheat flour =$0.7

cost per pound of the rye flour =$0.6